# Detecting and protecting entanglement through nonlocality, variational entanglement witness, and nonlocal measurements

### v1

@@ Abstract
We present an advanced method to enhance the detection and protection of quantum entanglement, a key concept in quantum mechanics for computing, communication, and beyond. Entanglement, where particles remain connected over distance, can be indicated by nonlocality, measurable through the Clauser-Horne-Shimony-Holt (CHSH) inequality. While violating the inequality confirms entanglement, entanglement can still exist without such violations. To overcome this limitation, we use the CHSH violation as an entanglement measure and introduce a variational entanglement witness for more complete detection. Moreover, we propose a nonlocal measurement framework to measure the expectation values in both the CHSH inequality and variational entanglement witness. These nonlocal measurements exploit the intrinsic correlations between entangled particles, providing a more reliable approach for detecting and maintaining entanglement. This paper significantly contributes to the practical application of quantum technologies, where detecting and maintaining entanglement are essential.
@@ I Introduction
Recent advancements in machine learning have introduced promising methods for detecting quantum entanglement. Neural networks and support vector machines have proven effective in classifying quantum states as either entangled or separable [33, 34, 35, 23, 36, 37, 38], detecting genuine multipartite entanglement [37], and developing entanglement witnesses [23, 20, 39]. So far, convolutional neural networks have been particularly useful for analyzing entanglement patterns [40, 36]. These methods highlight the expanding role of machine learning in enhancing the detection and analysis of quantum entanglement.
However, these detection methods rely on multiple measurements of a quantum state and local measurements on spatially separable subsystems, which cause the collapse of the global wavefunction of the entire system. Therefore, there is a need for detection methods that can identify entanglement without destroying it.
In this paper, we use the CHSH inequality violation as an indicator of entanglement. We also introduce a variational entanglement witness (VEW) to detect entanglement when the CHSH inequality alone is not enough. Moreover, we propose a nonlocal measurement framework to effectively measure the CHSH inequality and VEW, enabling both the detection and protection of entanglement.
In a bipartite system shared with Alice and Bob, we first theoretically examine the entanglement via the violation of the CHSH inequality. We then apply a VEW to independently detect entanglement by training a parameterized witness operator. Finally, we propose a nonlocal measurement framework to measure the expectation values in the CHSH and VEW. This framework improves the reliability of entanglement detection and protects it from wave function collapse. We also use superconducting chips to simulate the nonlocal measurement and assess the post-measurement state to confirm the preservation of entanglement. This work not only advances our fundamental understanding of quantum entanglement but also supports quantum applications like secure communication and complex computations.
@@ II.1 CHSH inequality and VEW
From this nonlocality, we can infer the entanglement in the quantum state $|\psi\rangle$. We confirm the entanglement using the PPT criterion and concurrence. Detailed calculations can be found in Appendix B. As depicted in Fig. 2(a), the maximum violation of inequality (2) corresponds to the highest level of entanglement, indicated by either minimum PPT or maximum concurrence. Thus, the violation of the CHSH inequality indicates the presence of entanglement. While the nonlocal behavior can indicate entanglement, however, entanglement does not always imply nonlocality, meaning nonlocality cannot fully detect entanglement.
@@
To find the most effective EW, here we propose a variational entanglement witness (VEW) for the first time, which optimizes $\mathcal{W}$ to better detect entanglement for a given quantum state. The variational scheme proceeds as follows. First, the entanglement witness $\mathcal{W}$ is parameterized by $\bm{\upalpha}$, ie., $\mathcal{W}(\bm{\upalpha})$. The expectation value of this witness is then used as the cost function
@@
In this work, we employ $S_{\rm CHSH}$ as the variational witness operator, thereby removing the requirement for prior knowledge of the quantum state and enabling direct measurement. The variational witness operator is defined by
@@
The optimization process bases on gradient-free methods using the COBYLA optimizer. See detailed calculation in Appendix C. The results, illustrated in Fig. 2(b), effectively demonstrate the entanglement of the quantum state $|\psi\rangle$. Specifically, at $\theta=0,\pi/2,$ and $\pi$, the expectation value $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle$ is zero, indicating zero in the PPT criterion and concurrence. At other points, $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle$ is negative, signifying entanglement. This observation holds for all $\phi$ except when $\phi=\pi/2$, because at this specific point, the expectation value $\langle XX\rangle=0$ and thus $\mathcal{C}(\bm{\upalpha})$ becomes a constant. According to the constrain in Eq. (7), this $\mathcal{C}(\bm{\upalpha})$ must be vanished.
We emphasize that both the CHSH inequality and VEW require nonlocal measurements of the expectation values $\langle ZZ\rangle$ and $\langle XX\rangle$. We will now present a framework for nonlocal measurements of $\langle ZZ\rangle$ and $\langle XX\rangle$ to confirm the CHSH inequality and VEW.
@@ II.2 Nonlocal measurement framework for measuring the CHSH inequality and VEW
To measure $\langle ZZ\rangle$, we apply the interaction $U_{1}$ between the system and M1 which are two CNOT (CX) gates as shown in Fig. 3(a). The measurement observables are represented by Kraus operators as
@@ II.3 Simulation on superconducting chip
where we used $2\theta$ in Eq. (22) to get $|\psi\rangle$. Similarly, to prepare $|\xi\rangle_{1}$, we apply a Hadamard gate onto $q_{2}$ followed by a CX gate $q_{2}q_{3}$, and to prepare $|\xi\rangle_{2}$, we apply a Hadamard gate onto $q_{4}$, X gate onto $q_{5}$, followed by a CX gate $q_{4}q_{5}$. The interaction $U_{1}$ consists of two CX gates, while $U_{2}$ consists of two inverted CX gates as shown in the figure. Measure $q_{2},q_{3}$ in the $Z$ basis gives the outcome for $\langle ZZ\rangle$. To get $\langle XX\rangle$, we apply Hadamard gates onto $q_{4}$ and $q_{5}$, and measure in the $Z$ basis.
@@ II.5 Mixed state case
See detailed calculation in Appendix F. Finally, we observe that the entanglement in this state remains unchanged from the initial state $\rho$, indicating that entanglement persists under nonlocal measurement.
@@ III Conclusion
We made significant progress in detecting and protecting quantum entanglement using nonlocality, variational entanglement witness (VEW), and nonlocal measurements. While traditional methods like violations of the CHSH inequality are effective, they do not cover all scenarios of entanglement detection. By introducing VEW, we addressed the limitations of these traditional methods, offered a more comprehensive approach to identifying entanglement. We also proposed a nonlocal measurement framework for measuring the CHSH operator and VEW.
Our findings emphasize the crucial role of nonlocal measurements in detecting and maintaining entanglement, which are essential for the functionality of quantum technologies. This work not only advances the understanding of quantum entanglement but also contributes to the practical development of more robust quantum computing, communication, and sensing systems.
@@ Appendix C Variation entanglement witness
In this section, we outline the optimization process for the variational entanglement witness. The cost function is defined as:
@@ Appendix D Nonlocal measurement
In this section, we derive detailed calculation of nonlocal measurement of $\langle ZZ\rangle$ and $\langle XX\rangle$.
@@ D.1 Nonlocal measurement of $\langle ZZ\rangle$
The measurement of $\langle ZZ\rangle$ is given though meter M1 initially prepare in $|\xi\rangle_{1}=\dfrac{1}{\sqrt{2}}\big{(}|\uparrow\uparrow\rangle+|\downarrow\downarrow\rangle\big{)}$. First, the interaction between the system and meter M1 is give through $U_{1}$, which are CX gates as
@@ D.2 Nonlocal measurement of $\langle XX\rangle$
Similar, to measure $\langle XX\rangle$, we use meter M2 with the initial state $|\xi\rangle_{2}=\dfrac{1}{\sqrt{2}}\big{(}|\circ\!\times\rangle+|\times\!\circ\rangle\big{)}$. The interaction $U_{2}$ is the invert CX gates
@@
where $q_{4}$ and $q_{5}$ are control qubits and $q_{0}$ and $q_{1}$ are target qubits. The action of $U_{2}$ on $|\xi\rangle_{2}$ gives
$U_{2}|\xi\rangle_{2}=\dfrac{1}{\sqrt{2}}\Big{(}IX\otimes|\circ\!\times\rangle+XI\otimes|\times\!\circ\rangle\Big{)}.$ (D.14)
After the interaction $U_{2}$, we apply the Hadamard gates ${\rm H}_{q_{4}}{\rm H}_{q_{5}}$ onto qubits $q_{4}$ and $q_{5}$ of the meter M2
@@
$=\dfrac{1}{2\sqrt{2}}\Big{[}\big{(}IX+XI\big{)}\otimes\big{(}|\circ\!\circ\rangle-|\times\!\times\rangle\big{)}+\big{(}IX-XI\big{)}\otimes\big{(}-|\circ\!\times\rangle+|\times\!\circ\rangle\big{)}\Big{]}$ (D.15)

### v2

@@ Abstract
We propose an innovative method to enhance the detection and protection of quantum entanglement, a cornerstone of quantum mechanics with applications in computing, communication, and beyond. While entanglement can be represented through nonlocal correlations detectable by the Clauser-Horne-Shimony-Holt (CHSH) inequality, this method does not fully capture all entangled states. To address this limitation, we introduce a variational entanglement witness (VEW) that optimizes the probabilities of detection and improves the efficiency of distinguishing between separable and entangled states. Additionally, we propose a novel nonlocal measurement framework that enables the assessment of both CHSH inequalities and the VEW while preserving the entanglement. Our approach enhances the reliability of entanglement detection while maintaining the entanglement of quantum states, offering significant advancements for quantum technologies.
@@ I Introduction
Recent advancements in machine learning have introduced promising methods for detecting quantum entanglement. Neural networks and support vector machines have proven effective in classifying quantum states as either entangled or separable [33, 34, 35, 23, 36, 37, 38, 39], detecting genuine multipartite entanglement [37], and developing entanglement witnesses [23, 20, 40, 39]. So far, convolutional neural networks have been particularly useful for analyzing entanglement patterns [41, 36]. These methods highlight the expanding role of machine learning in enhancing the detection and analysis of quantum entanglement.
However, these detection methods rely on multiple measurements of a quantum state and local measurements on spatially separable subsystems, which causes the collapse of the global wavefunction of the entire system. Therefore, there is a need for detection methods that can identify entanglement without destroying it.
In this paper, we introduce a variational entanglement witness (VEW) approach to detect entanglement when the CHSH inequality alone is not enough. While using these quantities for detecting entanglement is widespread, our approach provides an effective tool to confirm a quantum state entangled status. Optimizing VEW also improves the efficiency in distinguishing separable from entangled states. Moreover, we propose a nonlocal measurement framework to effectively measure the CHSH inequality and VEW, enabling both the detection and protection of entanglement.
Concretely, in a bipartite system shared with Alice and Bob, we first theoretically examine the entanglement via the violation of the CHSH inequality. We then apply a VEW to independently detect entanglement by training a parameterized witness operator. We extend the application of VEW to general pure and mixed states in two-dimensional systems and further to higher-dimensional systems. Finally, we propose a nonlocal measurement framework to measure the expectation values in the CHSH and VEW. This framework improves the reliability of entanglement detection and protects it from wave function collapse. We also use superconducting chips to simulate the nonlocal measurement and assess the post-measurement state to confirm the preservation of entanglement. This work not only advances our fundamental understanding of quantum entanglement but also supports quantum applications like secure communication and complex computations.
@@ II.1 CHSH inequality and VEW
From this nonlocality, we can infer the entanglement in the quantum state $|\psi\rangle$. We confirm the entanglement using the PPT criterion and concurrence. Detailed calculations can be found in Appendix B. As depicted in Fig. 2(a), the maximum violation of inequality (2) corresponds to the highest level of entanglement, indicated by either minimum PPT or maximum concurrence. Thus, the violation of the CHSH inequality indicates the presence of entanglement.
While the nonlocal behavior can indicate entanglement, however, entanglement does not always imply nonlocality, meaning nonlocality cannot fully detect entanglement.
@@
To find the most effective EW, here we propose a variational entanglement witness (VEW), which optimizes $\mathcal{W}$ to better detect entanglement for a given quantum state. The variational scheme proceeds as follows. First, the entanglement witness $\mathcal{W}$ is parameterized by $\bm{\upalpha}$, ie., $\mathcal{W}(\bm{\upalpha})$. The expectation value of this witness is then used as the cost function
@@
In principle, VEW can be constructed from any operator, including Bell’s operator, however it requires perfect correlations, which are impractical experimentally. Here, we employ $S_{\rm CHSH}$ as the variational witness operator, thereby removing the requirement for prior knowledge of the quantum state and enabling direct measurement. The variational witness operator is defined by
@@
The optimization process relies on gradient-free methods using the COBYLA optimizer. See detailed calculation in Appendix C.
II.1.1 Demonstration of VEW for a Bell state.
We first examine VEW for a Bell state given by $|\psi\rangle$ in Eq. (3). The results, illustrated in Fig. 2(b), effectively demonstrate the entanglement of the quantum state. At $\theta=0,\pi/2,$ and $\pi$, the expectation value $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle$ is zero, indicating zero in the PPT criterion and concurrence. At other points, $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle$ is negative, indicating the entanglement. Here, we achieve a 100% success rate, comparable to the performance of the machine learning approach [39]. For each point, we optimize VEW to obtain its minimum value. The minimum values are independent of one another, so the curves appear unsmooth. This observation holds for all $\phi$ except when $\phi=\pi/2$, because at this specific point, the expectation value $\langle XX\rangle=0$ and thus $\mathcal{C}(\bm{\upalpha})$ becomes a constant. According to the constraint in Eq. (7), this $\mathcal{C}(\bm{\upalpha})$ must vanish.
II.1.2 Demonstration of VEW for a general pure state.
For general cases, we examine a generic bipartite pure state represented by
$|\psi_{\rm gen}\rangle=a|00\rangle+b|01\rangle+c|10\rangle+d|11\rangle,$ (10)
where $a$, $b$, $c$, and $d$ are complex coefficients that satisfy the normalization condition $|a|^{2}+|b|^{2}+|c|^{2}+|d|^{2}=1$. The concurrence is calculated as
$C=2|ad-bc|.$ (11)
To evaluate numerically, we generate 100 random states, including 50 separable states ($C=0$) and 50 entangled states ($C>0$). VEW is then applied to classify the states as separable or entangled, with the results shown in Fig. 3. Refer to App. C for detailed information on the data generation process.
Figure 3(a) shows the set of random states, with blue dots representing separable states and red dots representing entangled states. The states are distributed randomly within an ellipse that illustrates the Hilbert space.
Figure 3(b) presents the classification results. The separable subspace is a convex subset that is nested within a larger convex set of all quantum states, while the entangled subspace is the region between them. Closed squares represent separable states (defined by $C=0$) and are placed in the separable subspace. Similarly, open squares represent entangled states and are assigned to the entangled subspace. Using VEW, these states are classified and assigned colors: blue means separable states with $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle\geq 0$ and red means entangled states with $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle<0$. The green dotted line indicates $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle=0$.
Correct classifications includes closed blue squares (separable states with positive VEW) and open red squares (entangled states with negative VEW). Misclassifications, on the other hand, appear as open blue squares (entangled states but positive VEW) and closed red squares (separable states but negative VEW). In this example of the random data set, VEW achieves 66% accuracy for separable states and 84% for entangled states. Moreover, states violating the CHSH inequality, marked by the smaller red dashed convex subset, are identified as entangled by both VEW and CHSH.
After all, we emphasize that both the CHSH inequality and VEW require nonlocal measurements of the expectation values $\langle ZZ\rangle$ and $\langle XX\rangle$. We will now present a framework for nonlocal measurements of $\langle ZZ\rangle$ and $\langle XX\rangle$ to confirm the CHSH inequality and VEW.
Figure 3: (a) A set of 100 random states, with 50 separable states shown as blue dots and 50 entangled states as red dots. The states are distributed within an ellipse representing the Hilbert space. (b) Classification results using VEW. The separable subspace is the blue ellipse, and the entangled subspace is the region between the red and blue ellipses. Closed squares indicate $C=0$ and opened squares are $C>0$. Blue colors indicate $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle\geq 0$ (separable), and red colors are $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle<0$ (entangled), separated by the green dotted line $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle=0$. The inset red dashed ellipse highlights the subspace where the CHSH inequality is violated.
@@ II.2 Nonlocal measurement framework for measuring the CHSH inequality and VEW
To measure $\langle ZZ\rangle$ (note that hereafter, we limit ourselves to the expectation values w.r.t the quantum state $|\psi\rangle$ in Eq. (3)), we apply the interaction $U_{1}$ between the system and M1 which are two CNOT (CX) gates as shown in Fig. 4(a). The measurement observables are represented by Kraus operators as
@@
Even though $\langle ZZ\rangle=1$ regardless of $\theta$ and $\phi$ however, this value is not always known beforehand in other scenarios. Therefore, using $U_{1}$ is necessary to transfer information from the system to the meter $M_{1}$, enabling the measurement outcomes to reveal information about the system.
@@ II.3 Simulation on superconducting chip
where we used $2\theta$ in Eq. (24) to get $|\psi\rangle$. Similarly, to prepare $|\xi\rangle_{1}$, we apply a Hadamard gate onto $q_{2}$ followed by a CX gate $q_{2}q_{3}$, and to prepare $|\xi\rangle_{2}$, we apply a Hadamard gate onto $q_{4}$, X gate onto $q_{5}$, followed by a CX gate $q_{4}q_{5}$. The interaction $U_{1}$ consists of two CX gates, while $U_{2}$ consists of two inverted CX gates as shown in the figure. Measure $q_{2},q_{3}$ in the $Z$ basis gives the outcome for $\langle ZZ\rangle$. To get $\langle XX\rangle$, we apply Hadamard gates onto $q_{4}$ and $q_{5}$, and measure on the $Z$ basis.
@@ II.5 Mixed state case
See detailed calculation in Appendix F. Finally, we observe that the degree of entanglement in this state remains unchanged from the initial state $\rho$, indicating that entanglement persists under nonlocal measurement.
We next extend the use of VEW to mixed states beyond the Werner state by considering the Bell-diagonal state, a mixture of the four Bell states
$\rho=\sum_{i=1}^{4}\lambda_{i}|\Lambda_{i}\rangle\langle\Lambda_{i}|,$ (38)
where $\lambda_{i}$ are eigenvalues obeying $\sum\lambda_{i}=1$, and $|\Lambda_{i}\rangle$ are the Bell states defined in Eqs. (30, 31). The entanglement of the state is determined by the largest eigenvalue $\lambda_{i}$: the state is separable if $\mathrm{max}(\lambda_{i})\leq 0.5$, and entangled otherwise.
Similar to the pure state case in Fig. 3, we numerically generate 100 random states, evenly split between 50 separable and 50 entangled, by controlling ${\rm max}(\lambda_{i})$ as illustrated in Fig. 7(a). VEW is then applied for classification, with the results shown in Fig. 7(b). VEW achieves classification accuracies of 66% for separable states and 70% for entangled states. In this example dataset, none of the entangled states violate the CHSH inequality. See also Fig. 9(b).
II.6 Demonstration of VEW for high-dimension systems
We have discussed various applications of the proposed VEW for two-dimensional bipartite systems with pure and mixed states. Now, we apply it to high-dimensional bipartite systems.
Figure 7: Classification of mixed states using VEW, similar to the pure states case in Fig. 3. (a) States are generated with ${\rm max}(\lambda_{i})\leq 0.5$ for separable states and ${\rm max}(\lambda_{i})>0.5$ for entangled states. (b) Classification results using VEW.
We consider a bipartite system of two qudit subsystems in the Hilbert space $H=\mathbb{C}_{A}^{d}\otimes\mathbb{C}_{B}^{d}$, with canonical bases $\{|i,i\rangle\equiv|i_{A}\rangle\otimes|i_{B}\rangle\}_{i=0}^{d-1}$, where $d$ is the dimension of each qudit. The Bell inequality generalizes to the Collins-Gisin-Linden-Massar-Popescu (CGLMP) inequalities for higher-dimensional systems [49]. We define the CGLMP operator $S_{\rm CGLMP}$ as
$S_{\rm CGLMP}=(X_{d}+Z_{d})\otimes P_{d}+(X_{d}-Z_{d})\otimes Q_{d},$ (39)
where the shift and phase operators $X_{d}$ and $Z_{d}$ are
$X_{kl}=\begin{cases}1&\text{if }l=(k+1)\mod d,\\ 0&\text{otherwise},\end{cases}$ (40)
$Z_{kl}=\begin{cases}\exp\left(\frac{2\pi i}{d}\cdot k\right)&\text{if }k=l,\\ 0&\text{otherwise},\end{cases}$ (41)
and $P_{d}=-(Z_{d}+X_{d})/\sqrt{2},Q_{d}=(Z_{d}-X_{d})/\sqrt{2}$.
The common used EW $\mathcal{W}_{d}$ is derived from $S_{\rm CGLMP}$ as
$\mathcal{W}_{d}=2\textbf{I}_{d}-S_{\rm CGLMP},$ (42)
where $\textbf{I}_{d}$ is the identity operator in $d$-dimensional space. We also define the VEW as
$\mathcal{W}_{d}(\bm{\upalpha})=-\sqrt{2}\left(\upalpha_{1}Z_{d}\otimes Z_{d}+\upalpha_{2}X_{d}\otimes X_{d}\right),$ (43)
and the cost function $\mathcal{C}_{d}(\bm{\upalpha})=\langle\mathcal{W}_{d}(\bm{\upalpha})\rangle$. The optimization problem is given by
[…]
Figure 8 shows the numerical results for several quantum states, including the separable state $|\psi\rangle=|0,0\rangle$ and two entangled states: $|\psi\rangle=\sum_{i}|i,i\rangle/\sqrt{d}$ (maximum entangled state) and $|\psi\rangle=(|0,0\rangle+|d-1,d-1\rangle)/\sqrt{2}$ (partial entangled state). We compare the performance of EW $\mathcal{W}_{d}$ (solid curves) and VEW $\mathcal{W}_{d}(\bm{\upalpha}^{*})$ (dotted curves). Although the EW does not distinguish between entangled and separable states (all $\mathcal{W}_{d}$ are negative), the VEW successfully classifies them ($\mathcal{W}_{d}(\bm{\upalpha}^{*})=0$ for separable state and $\mathcal{W}_{d}(\bm{\upalpha}^{*})<0$ for entangled states), demonstrating its advantage in detecting entanglement in high-dimensional systems.
Figure 8: Classification of quantum states in high-dimensional systems using EW $\mathcal{W}_{d}$ (solid curves) and VEW $\mathcal{W}_{d}(\bm{\upalpha}^{*})$ (dotted curves): EW fails to distinguish between entangled and separable states, whereas VEW successfully classifies them.
@@ III Conclusion
We made significant progress in detecting and protecting quantum entanglement using nonlocality, variational entanglement witness (VEW), and nonlocal measurements. While traditional methods like violations of the CHSH inequality are effective, they do not cover all scenarios of entanglement detection. By introducing VEW, we addressed the limitations of these traditional methods, and offered a more comprehensive approach to identifying entanglement. We also proposed a nonlocal measurement framework for measuring the CHSH operator and VEW. Our findings emphasize the crucial role of nonlocal measurements in detecting and maintaining entanglement, which are essential for the functionality of quantum technologies. This work not only advances the understanding of quantum entanglement but also contributes to the practical development of more robust quantum computing, communication, and sensing systems.
@@ Appendix C Variation entanglement witness
In this section, we outline the optimization process for the variational entanglement witness. The cost function is defined as
@@
Additional data for general random pure and mixed states. For general pure states, 100 random states were generated according to Eq. (10), consisting of 50 separable states ($C=0$) and 50 entangled states ($C>0$). The PPT, concurrence, and $\langle S_{\rm CHSH}\rangle$ values were analyzed, as shown in Fig. 9(a). Similarly, Fig. 9(b) presents results for 100 random mixed states, generated from Eq. (38), with 50 states satisfying $\text{max}(\lambda_{i})\leq 0.5$ and 50 states satisfying $\text{max}(\lambda_{i})>0.5$. These data are used to examine VEW, as detailed in Figs. 3 and 7 of the main text.
Figure 9: (a) Generating of 100 random pure states, with 50 states having $C=0$ and 50 states having $C>0$. For each state, the PPT, concurrence, and $\langle S_{\rm CHSH}\rangle$ values are displayed. (b) A similar process was applied to mixed states, with 50 states having ${\rm max}(\lambda_{i})\leq 0.5$ and 50 states having ${\rm max}(\lambda_{i})>0.5$ These data are used to examine VEW, as shown in Figs. (3, 7) of the main text.
@@ Appendix D Nonlocal measurement
In this section, we derive detailed calculations of nonlocal measurement of $\langle ZZ\rangle$ and $\langle XX\rangle$.
@@ D.1 Nonlocal measurement of $\langle ZZ\rangle$
The measurement of $\langle ZZ\rangle$ is given through meter M1 is initially prepared in $|\xi\rangle_{1}=\dfrac{1}{\sqrt{2}}\big{(}|\uparrow\uparrow\rangle+|\downarrow\downarrow\rangle\big{)}$. First, the interaction between the system and the meter M1 is given through $U_{1}$, which are CX gates as
@@ D.2 Nonlocal measurement of $\langle XX\rangle$
Similar, to measure $\langle XX\rangle$, we use meter M2 with the initial state $|\xi\rangle_{2}=\dfrac{1}{\sqrt{2}}\big{(}|\circ\!\times\rangle+|\times\!\circ\rangle\big{)}$. The interaction $U_{2}$ is the inverted CX gates
@@
where $q_{4}$ and $q_{5}$ are control qubits and $q_{0}$ and $q_{1}$ are target qubits. The action of $U_{2}$ on $I_{q_{0}}\otimes I_{q_{1}}\otimes|\xi\rangle_{2}$ gives
$U_{2}|\xi\rangle_{2}=\dfrac{1}{\sqrt{2}}\Big{(}IX\otimes|\circ\!\times\rangle+XI\otimes|\times\!\circ\rangle\Big{)},$ (D.14)
where, again, we omit $I_{q_{0}}\otimes I_{q_{1}}$ for short. After the interaction $U_{2}$, we apply the Hadamard gates ${\rm H}_{q_{4}}{\rm H}_{q_{5}}$ onto qubits $q_{4}$ and $q_{5}$ of the meter M2
@@
$=\dfrac{1}{2\sqrt{2}}\Big{[}\big{(}IX+XI\big{)}\otimes\big{(}|\circ\!\circ\rangle-|\times\!\times\rangle\big{)}+\big{(}IX-XI\big{)}\otimes\big{(}-|\circ\!\times\rangle+|\times\!\circ\rangle\big{)}\Big{]},$ (D.15)